Which of the following values comes closest to the net present value of a project that requires an initial investment of $250 and produces cash flows of $60 per year for 10 consecutive years beginning at the end of year 5 (the cash flows go from the end of year 5 through the end of year 14)? The required rate of return is 13%?

a.

$1.81

b.

($17.23)

c.

($64.70)

d.

($34.55)

e.

($50.32)
Answer #1

Ε F 017 fx =SUM(D2:016) B 1 Year Cash flows Pv@13% Present Value 2 0 $ (250.00) 1.0000 $ 250.00) 11 $ 0.8850 $ 2 $ - 0.7831 $

foc =SUM(D2:016) Cash flows Pv@13% 1 Year 20 Nm+mod =C2/1.13 =C3/1.13 =C4/1.13 =C5/1.13 =C6/1.13 =C7/1.13 =C8/1.13 =C9/1.13 =